# From Observation to Orientation: an Adaptive Integer Programming Approach to Intervention Design

### v1

@@ Abstract
Using both observational and experimental data, a causal discovery process can identify the causal relationships between variables. A unique adaptive intervention design paradigm is presented in this work, where causal directed acyclic graphs (DAGs) are for effectively recovered with practical budgetary considerations. In order to choose treatments that optimize information gain under these considerations, an iterative integer programming (IP) approach is proposed, which drastically reduces the number of experiments required. Simulations over a broad range of graph sizes and edge densities are used to assess the effectiveness of the suggested approach. Results show that the proposed adaptive IP approach achieves full causal graph recovery with fewer intervention iterations and variable manipulations than random intervention baselines, and it is also flexible enough to accommodate a variety of practical constraints.
@@ 1 Introduction
An intervention in a causal system amounts to exogenously setting (or forcing) one or more variables to specific values, effectively breaking their dependence on any natural parents in the graph [1]. Interventions offer an avenue to break the ambiguities present in essential graphs by “cutting” certain edges or revealing causal directions [2, 11]. In practical terms, interventions can be classified along several dimensions:
  • Single-variable vs. Multi-variable: Traditional designs like randomized controlled trials often intervene on a single “treatment” variable. More recent work considers manipulating multiple variables simultaneously. This can greatly accelerate the identification of causal edges but may be more expensive or technically challenging [5, 12]. More critically, multi-variable interventions are often necessary to uncover interaction effects, where the causal influence of one variable depends on the state of another. For example, intervening on multiple genes simultaneously in a biological pathway can reveal such complex interactions. Consider discovering how variables $A$ and $B$ affect a child variable $C$. If there is an interaction (e.g., the effect of $A$ on $C$ changes depending on the level of $B$), intervening only on $A$ (or only on $B$) might only reveal the effect at the current or average level of the non-intervened variable. However, by simultaneously intervening on both $A$ and $B$ (e.g., setting them to combinations like high/low levels in a factorial manner), one can directly compare outcomes to quantify the interaction, a relationship that single-variable interventions might obscure or cannot directly measure [13]. While graph structure (edge existence) might eventually be identifiable using only single interventions (albeit potentially inefficiently), understanding these functional interactions often requires manipulating causes together.
  • Perfect (hard) vs. Imperfect (soft): A perfect intervention fixes a variable to a given value, removing the influence of all its original parents. A gene knockout experiment, where a gene is completely disabled, is an example of a perfect intervention. In contrast, an imperfect or soft intervention modifies the variable’s distribution without necessarily erasing parental influences [14, 15]. For instance, using RNA interference to reduce the expression of a gene, rather than completely knocking it out, would be a soft intervention.
@@
Conventionally, experimental design in statistics has focused on how to best assign values (or treatment levels) within a fixed set of experimental factors [16, 13]. By contrast, in large-scale causal discovery, one must also decide which variables to intervene upon, often adaptively based on partial knowledge of the causal graph [6, 4, 8]. This introduces computational challenges, as the space of possible interventions grows exponentially with the number of variables. Broadly, two categories of intervention designs arise:
  • Fixed (passive) designs: The set of experiments is determined in advance, without considering any data or results from earlier experiments [5, 11, 17]. Although simpler, such designs can be highly inefficient, especially when the number of variables is large, as they do not leverage information gained during the discovery process.
  • Adaptive (active) designs: Each intervention is selected after observing the data from previous interventions, which refines one’s knowledge of the causal structure [6, 7]. These methods can dramatically reduce the total number of experiments by focusing on the remaining ambiguities in the partially learned graph.
@@
In this paper, we present an iterative method for intervention design framework towards adaptive causal structure discovery. In this paper, we focus on hard interventions, where each intervened variable is forcibly set, independently of its other causes. We present a novel integer programming (IP) framework that is able to
@@
    Select Variables Under Practical Constraints: We enforce a budgetary limit and a cap on the number of variables manipulated at once per intervention experiment. Our IP model determines which subset of variables to intervene on during each round of intervention experiment with the goal of maximizing the information gain. The gain is defined as the number of edges currently in an unknown status that could be updated after the experiment.
@@
    Handle Nonlinear Costs and Interaction Effects: We allow for the possibility that the cost of intervening on multiple variables simultaneously can be superadditive, i.e., significantly more expensive than simply adding the individual intervention costs.
  4. 4.
    Empirically Evaluate Efficiency: Through simulations on synthetic graphs of varying density and size, we show that our adaptive IP strategy outperforms the simple random intervention baseline, often requiring fewer rounds to fully identify the causal DAG.
Overall, our framework provides a flexible and efficient way to navigate the space of intervention experiments for causal discovery. The remainder of this paper is organized as follows. Section 2 provides relevant work. Section 3 formally states our problem setup and introduces the partially known graph (PKG) representation of a causal DAG. Section 4 details our proposed IP approach, including extensions for complex cost structures and integration with Meek’s rules. We then report on our experimental simulations and findings in Section 6, before concluding with a discussion of future directions.
@@ 2 Related Work
A diverse range of approaches have been explored for active learning in causal discovery. Decision-theoretic frameworks were employed by [19], incorporating intervention costs into their utility function and considering modular experiments where a single variable is targeted per intervention. They used an adaptive strategy. [20] focused on establishing theoretical bounds for identifying causal relations, considering scenarios where any number of variables could be simultaneously intervened on, However, they did not explicitly model intervention costs or provide a specific algorithm for intervention selection.[21] developed algorithms for both fixed and adaptive intervention design, aiming to minimize the number of manipulated variables, thus implicitly limiting intervention set size but without cost or budget considerations. [8] proposed the OPTINTER algorithm, incorporating limits on intervention set size through a parameter and employing an adaptive strategy. Their paper did not explicitly model intervention costs but aimed to minimize the number of experiments. Using a fixed strategy, [22] introduced greedy algorithms, OPTSINGLE, which employs an adaptive strategy and limits interventions to single vertices, and another algorithm, OPTUNB, which allows interventions on multiple variables simultaneously without a specified limit. The focus was on minimizing the number of interventions rather than the cost associated with each intervention. By drawing parallels to graph-theoretic concepts, [23] employed a fixed design strategy and implicitly addressed intervention costs by considering an objective of minimizing the number of intervenable variables in addition to the number of interventions. They provided algorithms to achieve this, considering worst-case (complete graph) scenarios which have limited practicality. [24] proposed theoretical limits to using randomized adaptive strategies and argued that they improve upon existing deterministic methods, but they did not incorporate limits on the number of variables that could be intervened on per intervention or model in their derived bound. Moreover, they do not provide algorithms to attain those limits. [25] considered both fixed and adaptive strategies, explicitly incorporating limits on intervention set size while aiming to minimize the number of interventions. While their adaptive algorithms meet the derived bounds, they do not consider intervention costs. Both [26] and [27] employ a fixed strategy and explicitly modeled intervention costs, with each variable having an associated cost. They formulated the problem of learning a causal graph as an integer program (IP), aiming to design a set of interventions with minimum total cost that can uniquely identify any causal graph with a given skeleton. The IP model was used to develop algorithms for specific graph structures, such as trees or clique trees. However, the cost structure is assumed to be additive. [28] employed a fixed strategy focusing on single-variable interventions and the number of interventions that can be performed. They did not explicitly consider intervention costs. [29] establish a new universal lower bound on the number of single-node interventions required to fully orient an essential graph. The paper did not propose any new algorithm but rather focused on establishing theoretical bounds. The proposed lower bound applies to any algorithm, whether adaptive or fixed. Finally, [17] allowed for explicit modeling of intervention (possibly nonlinear) costs and limits on intervention set size. They formulated a flexible IP model to determine the minimal set of interventions required for causal identifiability under a variety of practical scenarios. However, their approach was a fixed design strategy.
3 Problem Statement
@@
We address the adaptive design of interventions for causal structure learning, where, based on the current knowledge of causal structure, each intervention aims to reveal as much additional causal structure information as possible.
Partially Known Graph (PKG).
@@
After each intervention (and subsequent update rules), the PKG is revised based on newly identified adjacancies and orientations. We then apply Meek’s rules to further propagate orientations that do not require additional interventions. This PKG formalism allows us to encode a variety of background knowledge about the causal structure - including known or forbidden edges - throughout the discovery process. Note that the well-studied essential graph is a special case of the PKG that only involves $E_{\text{Adjacent}}$ and $E_{\text{Known}}$.
@@
Assumptions.
  1. 1.
  2. 2.
  3. 3.
Under these assumptions, our objective is to adaptively choose interventions that maximize the information gained about the true causal DAG, $G^{*}$, subject to some constraints such as a fixed intervention budget and/or a limit on how many variables can be intervened on in each experiment. After each intervention experiment, the PKG is updated, and Meek’s rules are applied to logically propagate further orientations.
4 Our Contribution
Building upon the previous work of [24, 26, 27, 17], we propose a novel iterative method to reduce uncertainty in a PKG by formulating an IP model augmented by an application of Meek’s rules. This approach is implemented as desrcibed in Algorithm 1.
Algorithm 1 Iterative Orientation via Adaptive IP and Meek’s Rules
  Input: A PKG with edges partitioned into: ${E}_{\text{Known}}$ (oriented edges), ${E}_{\text{Unknown}}$ (undecided), ${E}_{\text{semi}}$ (semi-directed), ${E}_{\text{Adjacent}}$ (undirected).
  Goal: Fully orient all edges into ${E}_{\text{Known}}$.
  repeat
     Run Adaptive_IP on PKG to propose orientations:
[…]
     Apply Meek’s Rules to orient additional edges:
      $\textsc{MeeksRules}(\text{{PKG}})\rightarrow\text{{PKG}}^{\prime}$.
     Replace PKG with the updated PKG.
  until all edges are placed into ${E}_{\text{Known}}$
  Output: Fully oriented graph in ${E}_{\text{Known}}$.
@@ 4.1 IP Model Formulation
In this section, the Adaptive_IP model is formulated, which aims to maximize the number of oriented edges at each iteration for a given PKG.
@@ 4.1.1 Notation
Maximum allowed number of variables that can be intervened on simultanenously.
@@ 4.1.3 Constraints
The budget constraint (Eq. (2)) restricts the combined cost of interventions and observations to remain within the allowed budget $B$, accounting for the intervention cost ($CI_{i}$) and observation cost ($CO_{i}$) for each variable. The next constraint (Eq. (3)) ensures that at most $k_{\text{max}}$ variables can be intervened on simultaneously, limiting the number of active manipulations per intervention. Orientation tests (Eqs. (4) and (5)) are defined such that an edge $(i,j)$ can only be subject to an orientation test if the source vertex $i$ is intervened on ($X_{i}=1$) and the target vertex $j$ is not ($X_{j}=0$). Similarly, adjacency tests (Eqs. (6) and (7)) require that neither vertex of an edge is intervened on ($X_{i}=0$ and $X_{j}=0$) for the test to apply. The update constraints (Eqs. (8), (9), and (10)) ensure that an edge in $E_{\text{Unknown}},E_{\text{Semi-directed}},$ or $E_{\text{Adjacent}}$ can only be updated if it is subject to the relevant orientation or adjacency tests. Finally, the binary constraint (Eq. (11)) forces all decision variables ($X_{i},O_{ij},A_{ij},IDU_{ij},IDS_{ij},IDA_{ij}$) to be binary.
4.2 IP Model Extensions
4.2.1 Cutting Planes
To strengthen the formulation of the Adaptive_IP model and eliminate infeasible solutions, the following cutting planes are introduced:
Extended Mutual Exclusion:
These two constraints can also be combined. This is enforced with the constraint:
$O_{ij}+O_{ji}+A_{ij}\leq 1\quad\forall(i,j)\in E_{\text{Unknown}}\cup E_{\text{Semi-directed}}$
4.2.2 Non-linear Cost Structures
Example (Two Variables):
Let $Y_{12}$ be a binary variable indicating whether both $V_{1}$ and $V_{2}$ are intervened on simultaneously ($Y_{12}=1$) or not ($Y_{12}=0$). Define a "delta cost," $C_{12}$, representing the extra cost incurred when both variables are jointly intervened on beyond their individual sums. In this example, $C_{12}=CI_{1,2}-CI_{1}-CI_{2}=10-1-1=8$.
The budget constraint becomes:
$CI_{1}\cdot X_{1}+CI_{2}\cdot X_{2}+C_{12}\cdot Y_{12}$ $\leq B$ (12)
$Y_{12}$ $\leq X_{1}$ (13)
$Y_{12}$ $\leq X_{2}$ (14)
$Y_{12}$ $\geq X_{1}+X_{2}-1$ (15)
These constraints ensure $Y_{12}=1$ if and only if both $X_{1}=1$ and $X_{2}=1$.
Generalization to Multiple Variables:
$Y_{S}$ $\leq X_{i_{j}},\quad\forall j$ (16)
$Y_{S}$ $\geq\sum_{j=1}^{k}X_{i_{j}}-(k-1)$ (17)
The budget constraint then incorporates these interactions:
$\sum_{i}CI_{i}\cdot X_{i}+\sum_{S}C_{S}\cdot Y_{S}$ $\leq B$ (18)
Physical Limitations and Prohibitive Costs:
4.2.3 Alternative Objective Functions
1. Average Gain over MEC (Ghassami et al. [28]):
2. Worst-Case Gain (Minimax Objective):
$\text{maximize}_{I}\quad\min_{G_{i}\in\text{MEC}(G^{*})}|R(I,G_{i})|$
3. Weighted Objective (Prioritizing Informative Edges):
How weights are assigned is crucial and heuristic. Examples include:
4. Targeted Discovery Objective:
5. Cost-Effectiveness (Approximation):
[…]
Here, $\lambda\geq 0$ is a penalty parameter balancing information gain (potentially weighted) against intervention cost. This encourages finding solutions that are not just informative but also relatively cheap, though it requires tuning $\lambda$. Note that the budget constraint (Eq. 2) still applies.
4.2.4 Batch Interventions for Parallel Execution
The objective would typically aim to maximize the totalinformation gained across the batch:
Constraints need modification:
  • Budget per Experiment: A budget $B_{b}$ could apply to each experiment $b$.
    $\sum_{i\in V}(CI_{i}X_{i,b}+CO_{i}(1-X_{i,b}))\leq B_{b}\quad\forall b$
  • Intervention Limit per Experiment:
    $\sum_{i\in\mathcal{X}}X_{i,b}\leq k_{\text{max}}\quad\forall b$
    $\sum_{b=1}^{k_{\text{batch}}}\sum_{i\in V}CI_{i}X_{i,b}\leq B_{\text{total}}$
    $IDU_{ij,b}\leq O_{ij,b}+O_{ji,b}+A_{ij,b},\quad\dots\quad\forall b$
4.2.5 Dynamic Constraints
4.2.6 Refined Cost Modeling for Observations
The budget constraint (Eq. 2) assumes an observation cost $CO_{i}$ is incurred for every variable $i$ that is not intervened on ($X_{i}=0$). This might be an overestimation. A more refined model could link observation costs directly to the tests being performed. For example, $CO_{i}$ is incurred only if $i$ is the target of an orientation test ($\exists j:O_{ji}=1$) or involved in an adjacency test ($\exists j:A_{ij}=1\text{ or }A_{ji}=1$).
Let $Obs_{i}$ be a binary variable, $Obs_{i}=1$ if node $i$ needs to be observed.
$Obs_{i}\geq O_{ji}\quad\forall j$
$Obs_{i}\geq A_{ik}\quad\forall k$
$Obs_{i}\geq A_{ki}\quad\forall k$
The budget constraint would then use $\sum CO_{i}\cdot Obs_{i}$ for the observation costs, potentially reducing the estimated cost compared to the simpler $(1-X_{i})$ formulation. This adds complexity but might allow for more informative interventions within the same budget if observation costs are significant.
@@ 4.3 Transition Rules for Edge Sets
$\mathbf{E_{\text{Unknown}}}$ $O_{ij}=1$, presence. $O_{ij}=1$, absence. (Add in Reverse direction). $A_{ij}=1$, presence.
$\mathbf{E_{\text{Semi-directed}}}$ $O_{ij}=1$ or $A_{ij}=1$, presence.
$\mathbf{E_{\text{Adjacent}}}$ $O_{ij}=1$ or $O_{ji}=1$, presence.
Table 1: Conditions and actions for edge transitions between edge sets. Presence/absence refers to if the intervention reveals the edge to be present/absent, respectively. This is done through conditional independence tests.
4.3.1 Transition Rules for Edge Sets
@@
If the edge $(i,j)$ is subject to an orientation test ($O_{ij}=1$) and the intervention reveals the edge to be present, then the edge is added to $E_{\text{Known}}$ in the tested direction. Otherwise, if the edge is absent, it is added to $E_{\text{Semi-directed}}$ in the reverse direction $(j\to i)$. If the edge $(i,j)$ is subject to an adjacency test ($A_{ij}=1$) and the intervention reveals that its present, the edge is moved to $E_{\text{Adjacent}}$. Once moved into another edge set, the edge is removed from $E_{\text{Unknown}}$. Also, if the edge is subject to an adjacency test and the intervention reveals that the edge is absent, it is removed from $E_{\text{Unknown}}$.
@@
If the edge $(i,j)$ is subject to an orientation test ($O_{ij}=1$) or adjacency test ($A_{ij}=1$) and the intervention reveals the edge to be present, it is moved to $E_{\text{Known}}$ with its resolved direction. The edge is then removed from $E_{\text{Adjacent}}$. and likewise if the intervention reveals the edge to be absent.

### v3

@@ Abstract
Using both observational and experimental data, a causal discovery process can identify the causal relationships between variables. A unique adaptive intervention design paradigm is presented in this work, with the objective of learning causal directed acyclic graphs (DAGs) under practical considerations and common assumptions. An iterative integer programming (IP) framework is proposed, which drastically reduces the number of interventions required. Simulations over a broad range of graph topologies are used to assess the effectiveness of the proposed approach. Results show that the proposed adaptive IP approach achieves full causal graph recovery with fewer interventions and variable manipulations than random intervention baselines, while remaining flexible to practical constraints.
@@ 1 Introduction
Interventions in a causal system exogenously set variable values, breaking dependencies [1] to resolve causal ambiguities [2, 11]. Key types include:
  • Single-variable vs. Multi-variable: Traditional single-variable interventions contrast with multi-variable approaches that accelerate discovery [5, 12] and are necessary for capturing interaction effects [13].
  • Perfect (hard) vs. Imperfect (soft): Perfect interventions fix values, removing parent influences (e.g., gene knockout). Imperfect interventions modify distributions while retaining some parental influence (e.g., gene knockdown) [14, 15].
@@
Conventionally, experimental design in statistics has focused on how best to assign treatment levels within a predetermined set of experimental factors [16, 13]. Large-scale causal discovery poses an additional challenge: one must also decide which variables to intervene on, and that choice typically depends on the partial knowledge that has accumulated about the causal graph [6, 4, 8]. Because the number of possible intervention subsets grows exponentially with the number of variables, this decision problem quickly becomes computationally demanding. Two broad approaches are employed. In a fixed or passive design, the entire sequence of experiments is planned in advance and executed exactly as specified, without regard to any information revealed along the way [5, 11, 17]. While conceptually simple, a fixed schedule can be highly inefficient for large systems because it ignores opportunities to redirect effort toward the remaining uncertainties in the graph. By contrast, an adaptive or active design selects each new intervention only after the data from previous rounds have been analysed [6, 7]. This feedback loop allows the experimenter to focus on the edges and orientations that are still ambiguous, and in many practical settings it dramatically reduces the total number of experiments required to recover the full causal structure.
@@
In this paper, we present an iterative method for intervention design inline with adaptive intervention design for causal discovery. In this paper, we assume hard interventions, where each intervened variable is forcibly set, independently of its direct causes. We present a novel integer programming (IP) framework that can:
@@
    Select Informative Intervention Variables: Our IP model determines which subset of variables to intervene on during each round of intervention experiment with the goal of maximizing the knowledge gained about the graph.
@@
    Integrate Practical Limitations as Constraints The proposed IP is highly modular and offers great flexibility. It accommodates various practical considerations as constraints and can be easily extended to accommodate a number of scenarios.
The remainder of this paper is organized as follows. Section 2 provides relevant work. Section 3 formally states our problem setup and introduces the partially known graph (PKG) representation of a causal DAG. It also details our proposed adaptive IP approach, and integration with Meek’s rules. We then report on our experimental simulations and findings in Section 4, before concluding with a discussion of future directions.
@@ 2 Related Work
A diverse range of approaches have been explored for active learning in causal discovery. Decision-theoretic frameworks were employed by [19], incorporating intervention costs into their utility function and considering modular experiments where a single variable is targeted per intervention. They used an adaptive strategy. [20] focused on establishing theoretical bounds for identifying causal relations, considering scenarios where any number of variables could be simultaneously intervened on, However, they did not explicitly model intervention costs or provide a specific algorithm for intervention selection.[21] developed algorithms for both fixed and adaptive intervention design, aiming to minimize the number of manipulated variables, thus implicitly limiting intervention set size but without cost or budget considerations. [8] proposed the OPTINTER algorithm, incorporating limits on intervention set size through a parameter and employing an adaptive strategy. Their paper did not explicitly model intervention costs but aimed to minimize the number of experiments. Using a fixed strategy, [22] introduced greedy algorithms, OPTSINGLE, which employs an adaptive strategy and limits interventions to single vertices, and another algorithm, OPTUNB, which allows interventions on multiple variables simultaneously without a specified limit. The focus was on minimizing the number of interventions rather than the cost associated with each intervention. By drawing parallels to graph-theoretic concepts, [23] employed a fixed design strategy and implicitly addressed intervention costs by considering an objective of minimizing the number of intervenable variables in addition to the number of interventions. They provided algorithms to achieve this, considering worst-case (complete graph) scenarios which have limited practicality. [24] proposed theoretical limits to using randomized adaptive strategies and argued that they improve upon existing deterministic methods, but they did not incorporate limits on the number of variables that could be intervened on per intervention or model in their derived bound. Moreover, they do not provide algorithms to attain those limits. [25] considered both fixed and adaptive strategies, explicitly incorporating limits on intervention set size while aiming to minimize the number of interventions. While their adaptive algorithms meet the derived bounds, they do not consider intervention costs. Both [26] and [27] employ a fixed strategy and explicitly modeled intervention costs, with each variable having an associated cost. They formulated the problem of learning a causal graph as an integer program (IP), aiming to design a set of interventions with minimum total cost that can uniquely identify any causal graph with a given skeleton. The IP model was used to develop algorithms for specific graph structures, such as trees or clique trees. However, the cost structure is assumed to be additive. [28] employed a fixed strategy focusing on single-variable interventions and the number of interventions that can be performed. They did not explicitly consider intervention costs. [29] establish a new universal lower bound on the number of single-node interventions required to fully orient an essential graph. The paper did not propose any new algorithm but rather focused on establishing theoretical bounds. The proposed lower bound applies to any algorithm, whether adaptive or fixed. Finally, [17] allowed for explicit modeling of intervention (possibly nonlinear) costs and limits on intervention set size. They formulated a flexible IP model to determine the minimal set of interventions required for causal identifiability under a variety of practical scenarios. However, their approach was for a fixed design strategy.
3 Methodology
@@
We address the adaptive design of interventions for causal structure learning, where, based on the current knowledge of causal structure, each intervention aims to reveal as much additional causal structure knowledge as possible.
Assumptions.
  1. 1.
  2. 2.
  3. 3.
Under these assumptions, and building upon the previous work of [24, 26, 27, 17], we solve an IP model, Adaptive_IP, to adaptively choose the set of variables to intervene on. We choose the set which maximizes the knowledge learnt about the true causal DAG, $G^{*}$, subject to some constraints such as a fixed intervention budget and/or a limit on how many variables can be intervened on in each experiment. After each intervention, the PKG is updated, and Meek’s rules are applied to logically propagate further orientations. This approach is implemented as described in Algorithm 1. Proofs of correctness and finite convergence of this algorithm can be found in Appendix B.
Algorithm 1 Iterative Orientation via Adaptive IP and Meek’s Rules
  Input: A PKG with edges partitioned into: ${E}_{\text{Known}}$ (oriented edges), ${E}_{\text{Unknown}}$ (undecided), ${E}_{\text{semi-directed}}$ (semi-directed), ${E}_{\text{Adjacent}}$ (undirected).
  Goal: Fully orient all edges into ${E}_{\text{Known}}$.
  repeat
     Run Adaptive_IP on PKG to propose orientations:
[…]
     Apply Meek’s Rules to orient additional edges:
      $\textsc{MeeksRules}(\text{{PKG}})\rightarrow\text{{PKG}}^{\prime}$.
     Replace PKG with the updated PKG.
  until all edges are placed into ${E}_{\text{Known}}$
  Output: Fully oriented graph in ${E}_{\text{Known}}$.
Partially Known Graph.
@@
After each intervention, the PKG is revised based on newly identified adjacencies and orientations as shown in section 3.2.
We then apply Meek’s rules to further propagate orientations that do not require additional interventions. This PKG formalism allows us to encode a variety of background knowledge about the causal structure - including known or forbidden edges - throughout the discovery process. Note that the well-studied essential graph is a special case of the PKG that only involves $E_{\text{Adjacent}}$ and $E_{\text{Known}}$.
@@ 3.1 IP Model Formulation
We formulate the Adaptive_IP model as follows:
@@ 3.1.1 Notation
Maximum allowed number of variables that can be intervened in simultaneously.
@@ 3.1.3 Constraints
The budget constraint (Eq. (2)) restricts the combined cost of interventions and observations to remain within the allowed budget $B$, accounting for the intervention cost ($CI_{i}$) and observation cost ($CO_{i}$) for each variable. The next constraint (Eq. (3)) ensures that at most $k_{\text{max}}$ variables can be intervened on simultaneously, limiting the number of active manipulations per intervention. Orientation tests (Eqs. (4) and (5)) are defined such that an edge $(i,j)$ can only be subject to an orientation test if the source vertex $i$ is intervened on ($X_{i}=1$) and the target vertex $j$ is not ($X_{j}=0$). Similarly, adjacency tests (Eqs. (6) and (7)) require that neither vertex of an edge is intervened on ($X_{i}=0$ and $X_{j}=0$) for the test to apply. The update constraints (Eqs. (8), (9), and (10)) ensure that an edge in $E_{\text{Unknown}},E_{\text{Semi-directed}},$ or $E_{\text{Adjacent}}$ can only be updated if it is subject to the relevant orientation or adjacency tests. Finally, the binary constraint (Eq. (11)) forces all decision variables to be binary. We wish to maximize the sum of edge updates given by objective 1.
The model is quite modular and offers flexibility to accommodate various alternative objectives and constraints. We showcase a wide variety of model extensions in Appendix C.
@@ 3.2 Transition Rules for Edge Sets
Table 1: Conditions and actions for edge transitions between edge sets.
@@
$\mathbf{E_{\text{Unknown}}}$ $O_{ij}=1$ presence. $O_{ij}=1$ absence. $A_{ij}=1$ presence.
$\mathbf{E_{\text{Semi-directed}}}$ ($O_{ij}=1$ or $A_{ij}=1$) & presence.
$\mathbf{E_{\text{Adjacent}}}$ ($O_{ij}=1$ or $O_{ji}=1$) & presence.
@@
If the edge $(i,j)$ is subject to an orientation test ($O_{ij}=1$) and the intervention reveals the edge to be present, then the edge is added to $E_{\text{Known}}$ in the tested direction. Otherwise, if it is absent, it is added to $E_{\text{Semi-directed}}$ in the reverse direction $(j\to i)$. If the edge $(i,j)$ is subject to an adjacency test ($A_{ij}=1$) and the intervention reveals that its present, the edge is moved to $E_{\text{Adjacent}}$. Once moved into another edge set, the edge is removed from $E_{\text{Unknown}}$. Also, if the edge is subject to an adjacency test and the intervention reveals that the edge is absent, it is removed from $E_{\text{Unknown}}$.
@@
If the edge $(i,j)$ is subject to an orientation test or an adjacency test, and the intervention reveals the edge to be present, it is moved to $E_{\text{Known}}$ with its resolved direction. The edge is then removed from $E_{\text{Semi-directed}}$ and likewise if the intervention reveals it to be absent.